### v2

# On the distance between mean and geometric median in high dimensions

Abstract
The geometric median, a notion of “center” for multivariate distributions, has gained recent attention in robust statistics and machine learning. Although conceptually distinct from the mean (i.e., expectation), we demonstrate that both are very close in high dimensions when the dependence between the distribution components is suitably controlled. Concretely, we find an upper bound on the distance that vanishes with the dimension asymptotically, and derive a rate-matching first order expansion of the geometric median components. Simulations illustrate and confirm our results.
keywords:
spatial median, robust mean estimation
1 Introduction
The geometric median provides a robust alternative to the sample mean, mitigaging the impact of outliers. The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which in practice often is the learning target [^1] For special distributions like the Gaussian, geometric median and mean can coincide. . A method to balance robustness and bias is the median-of-means [8, 10] or “bucketing” approach [5].
In this paper, we take a step back and quantify the actual bias of the geometric median, focusing on high-dimensional distributions. We prove, when each distribution component only depends on a fixed number of neighboring components, that the maximum-norm distance between geometric median and mean vector is of order $\mathcal{O}(\tfrac{1}{p})$, where $p$ denotes the dimension. Furthermore, we show that this rate is sharp in general.
The results are discussed at the end of the paper. Proofs are deferred to the appendix.
2 Theoretical results
We consider random vectors $\mathbf{X}:=(X_{1},\ldots,X_{p})\in\mathbb{R}^{p}$ for $p\in\mathbb{N}$, where the sequence of component random variables $(X_{i})_{i\in\mathbb{N}}$ is from the class $\mathcal{D}\equiv\mathcal{D}(M,q,c,\sigma_{\min},C)$ introduced below. A geometric median of $\mathbf{X}$ is any vector $\mathbf{m}\in\mathbb{R}^{p}$ minimizing the expectation $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ for the Euclidean norm $\|\cdot\|$. By properties of the class $\mathcal{D}$, there exists a unique geometric median (Lemma A.1), which we denote by $\mathbf{m}_{p}$. We then study $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|$, where $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$.
Definition 2.1.
Let $M\geq 0$ be an integer, and let $c,\sigma_{\min}>0$, $q\geq 2$ and $C\geq 1$ be reals. The class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ is comprised of all sequences of absolutely continuous r.v. $(X_{i})_{i\in\mathbb{N}}$ that are $M$-dependent, with $\operatorname{E}\left[\,|X_{i}|^{q}\,\right]\leq C$, $\operatorname{E}\left[\,|X_{i}-\mu_{i}|^{q}\,\right]\leq C$, and $\operatorname{sd}\left[\,X_{i}\,\right]\geq\sigma_{\min}>0$, and for which the probability density of $X_{i}$ is bounded by $c<\infty$ for all $i\in\mathbb{N}$.
Recall that $M$-dependence requires independence of $(X_{i})_{i\leq k}$ and $(X_{i})_{i\geq k+M+1}$ for all $k\in\mathbb{N}$. Note that $|\mu_{i}|\leq C$ and $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ since $C\geq 1$.
We build on the following central identity, derived from $\nabla_{\mathbf{m}}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]|_{\mathbf{m}=\mathbf{m}_{p}}=0$ by interchanging derivative with expectation, and then adding $\pm\boldsymbol{\mu}$:
$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right],$ (1)
where all expectations exist when $p$ is large enough (Lemma A.2).
The intuition is now that, by $M$-dependence, $X_{i}-\mu_{i}$ and $\|\mathbf{X}-\mathbf{m}_{p}\|$ are nearly independent for $p$ large, and thus $m_{p,i}-\mu_{i}\approx\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\cdot 1=0$ by (1). Formalizing this approach requires controlling $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{-1}\,\right]$, or slight variations involving $\check{\mathbf{X}}$ introduced below, uniformly in the dimension $p$. To this end, we combine the two lemmas, where the first is an initial bound.
Lemma 2.2.
When $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$, then $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\sqrt{Cp}$.
Lemma 2.3.
Let $(Y_{i})_{i\in\mathbb{N}}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$ and $\mathbf{v}_{p}\in\mathbb{R}^{p}$ with $\|\mathbf{v}_{p}\|^{2}\leq C_{2}p$ for $p\in\mathbb{N}$ and some $C_{2}<\infty$. With $\mathbf{Y}:=(Y_{1},\ldots,Y_{p})$ we have
$\sup_{p\geq\lceil 4k\rceil(M+1)+1}\operatorname{E}\left[\,\left(\frac{\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}\,\right]}\right)^{-k}\,\right]=\beta(k,C_{2},\mathcal{D})<\infty,$ (2)
for any $k>0$ and constant $\beta$ depending only on $k,C_{2}$ and the $\mathcal{D}$-parameters.
Both Lemmas together with the Cauchy-Schwarz inequality applied to the second expectation in (1) lead to the following intermediate bound.
Lemma 2.4.
If $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, then for a constant $B_{1}$ only depending on the $\mathcal{D}$-parameters, $\sup_{p\in\mathbb{N};\;1\leq i\leq p}|\mu_{i}-m_{p,i}|\leq B_{1}$.
We can now derive the main results, which refine the Cauchy-Schwarz bound in the proof of Lemma 2.4. More precisely, we show that $\operatorname{E}\left[\,(X_{i}-\mu_{i})/\|\mathbf{X}-\mathbf{m}_{p}\|\,\right]$ is close to zero by inserting $0=\operatorname{E}\left[\,(X_{i}-\mu)/\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\,\right]$, where $\check{\mathbf{X}}$ is missing all components with indices in $i-M,\ldots,i+M$. Formally, using the exclusion set $\mathcal{E}(i):=\mathbb{N}\cap\{i-M,\ldots,i+M\}$, we define $\check{\mathbf{v}}:=(v_{i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}(i)}\in\mathbb{R}^{|\{1,\ldots,p\}\setminus\mathcal{E}(i)|}$ for any vector $\mathbf{v}=(v_{i})_{i=1}^{p}\in\mathbb{R}^{p}$. Thus, subtracting zero from the right hand side of (1), a bit of algebra yields that for $p\geq i+M$,
$m_{p,i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right].$ (3)
The first expectation is $\mathcal{O}(\sqrt{p})$ while the second is $\mathcal{O}(\tfrac{1}{p^{3/2}})$, leading to
Theorem 2.5.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$. Then, with $B_{3}$ depending only on the $\mathcal{D}$-parameters, $|m_{p,i}-\mu_{i}|\leq\frac{B_{3}}{p}$ for all $1\leq i\leq p\in\mathbb{N}$.
In particular, Theorem 2.5 implies $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\tfrac{B_{3}}{\sqrt{p}}$.
By $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\to 0$, we can identify the limit in probability of the (suitably scaled) expectation arguments in (3). A uniform integrability argument (requiring $q>3$) transfers the convergence to the expectations themselves and yields
Theorem 2.6.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>3$. Fix $i\in\mathbb{N}$. If $\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\to\bar{\sigma}^{2}\in\mathbb{R}$, then
$m_{p,i}=\mu_{i}-\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right]+o\left(\frac{1}{p}\right).$ (4)
Note that while $\sigma_{\min}^{2}\leq\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ always has a convergent subsequence, there may be multiple accumulation points.
3 Simulations
We simulate $\hat{\mathbf{m}}_{p}$ for a range of distributions to validate Theorems 2.5 and 2.6. The simulation details are discussed at the end of this section.
Figure 1: (Left) Euclidean distance $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$. (Right) Residuals from Theorem 2.6 for $i=1$; see equation (5). Black reference lines are $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ (left) and $\mathcal{O}(\tfrac{1}{p})$ (right).
We plot $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$, for which Theorem 2.5 predicts $\mathcal{O}(\tfrac{1}{\sqrt{p}})$, on the left of Figure 1. On the right, we plot the residual
$\left|\hat{m}_{p,1}-\mu_{1}+\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(1)}\operatorname{E}\left[\,(X_{1}-\mu_{1})(X_{j}-\mu_{j})^{2}\,\right]\right|,$ (5)
for which Theorem 2.6 predicts $o(\tfrac{1}{p})$. We consider four data-generating distributions $(X_{i})_{i\in\mathbb{N}}$: first, independent exponential random variables with rates uniformly sampled from $[0.75,1.25]$, second an MA(2) process $X_{i}=L_{i}+0.2L_{i-1}-0.1L_{i-2}$ with iid skew normal noise $(L_{i})_{i}$, third iid Pareto variables with shape parameter $3.1$ (meaning $\operatorname{E}\left[\,|X_{1}|^{k}\,\right]<\infty$ iff $k<3.1$), and finally independent Paretos with shapes uniformly sampled from $[7,8]$. Table 1 lists linear regression slopes to the simulation curves after $p\geq 100$.
The simulation study clearly supports the theory, although $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ approaches $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ rather slowly for the heavy-tailed Pareto distributions. To see why, note $m_{p,1}-\mu_{1}\approx\tfrac{-19.5}{p}+\tfrac{20.2}{p^{1.03}}$ for Pareto(3.1) from the component expansion plot. Since all geometric median components are equal in the iid setting, $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\approx\sqrt{p}\cdot|\tfrac{-19.5}{p}+\tfrac{20.2}{p^{1.03}}|$, which indeed fits the simulations results (Figure 2). The near cancellation in the absolute value slows the decay for moderate $p$.
[…]
4 Discussion
The bias of the geometric median was recently highlighted in [11], where a constant upper bound on the Euclidean bias for a large class of distributions is derived. We complement this result by showing the Euclidean bias to be $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ for a smaller class.
From an estimation viewpoint, a small distance to the mean is desirable. From a robustness viewpoint it raises concern, as the mean is not robust against, say, contaminating a portion of the observations. A differentiation clarifies this: while $M$-dependence is compatible with certain cellwise contamination where indeed the geometric median is not always robust [12], rowwise contaminated observations are typically not $M$-dependent and the geometric median is robust here [10, 13].
There is empirical evidence that the geometric median may similarly be close to the mean for sparsely interacting random fields [13].
Our results link to the classic paper by [1] showing that (sample) geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
Finally, we remark that the expansion in Theorem 2.6 always contains the term $\operatorname{E}\left[\,(X_{i}-\mu_{i})^{3}\,\right]$, i.e., the negative component skewness. This aligns with statistical intuition for the univariate median [6].
Acknowledgements This paper is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research. Further, this work has been funded by the German Federal Ministry of Education and Research and the Bavarian State Ministry for Science and the Arts. The authors of this work take full responsibility for its content.
Appendix A Appendix
A.1 Proofs
Lemma A.1.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ and $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$. Then, when $p\geq M+2$, there exists a unique geometric median of $\mathbf{X}$.
Proof.
For $\mathbf{m},\mathbf{m}^{\prime}\in\mathbb{R}^{p}$, we have that
$|\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]-\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}^{\prime}\|\,\right]|\leq\operatorname{E}\left[\,\|\mathbf{m}-\mathbf{m}^{\prime}\|\,\right]=\|\mathbf{m}-\mathbf{m}^{\prime}\|$
by the inverse triangle inequality, hence $\mathbf{m}\mapsto\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ is continuous. Also, $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ diverges to $\infty$ as $\|\mathbf{m}\|\to\infty$, since
$\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]\geq\operatorname{E}\left[\,|\|\mathbf{X}\|-\|\mathbf{m}\||\,\right]\geq\|\mathbf{m}\|-\operatorname{E}\left[\,\|\mathbf{X}\|\,\right].$
Since continuous functions admit minimizers on compact sets, a geometric median exists.
When $p\geq M+2$, then $X_{1}$ and $X_{M+2}$ are independent, which ensures the distribution of $\mathbf{X}$ does not concentrate on a one-dimensional linear subspace. By [9], the geometric median is unique. ∎
Lemma A.2.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Set $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$, denote $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ and let $\mathbf{m}_{p}$ be the geometric median of $\mathbf{X}$. Then, when $p\geq 2(M+1)+1$,
$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$
Proof.
Let $G(\mathbf{m}):=\|\mathbf{X}-\mathbf{m}\|$ for $\mathbf{m}\in\mathbb{R}^{p}$. We have
$|\partial_{i}G(\mathbf{m})|=\left|\frac{(m_{i}-X_{i})}{\|\mathbf{X}-\mathbf{m}\|}\right|\leq\frac{\|\mathbf{X}-\mathbf{m}\|_{1}}{\|\mathbf{X}-\mathbf{m}\|}\leq\sqrt{p}$
by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [14, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
$\nabla\operatorname{E}\left[\,G(\mathbf{m})\,\right]=\operatorname{E}\left[\,\frac{\mathbf{m}-\mathbf{X}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$
Since $\mathbf{m}_{p}$ optimizes $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$, we have
$0=\operatorname{E}\left[\,\frac{\mathbf{X}-\mathbf{m}_{p}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$ (6)
By Lemma 2.2 and 2.3 with $k=\tfrac{1}{2}$ (both proven separately later), we have for $p\geq 2(M+1)+1$ that
$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]\leq\frac{\beta(0.5,C,\mathcal{D})}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\leq\frac{\beta(0.5,C,\mathcal{D})}{\sigma_{\min}\sqrt{p}}<\infty$
by $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\geq\operatorname{E}\left[\,\|\mathbf{X}-\boldsymbol{\mu}\|^{2}\,\right]$ due to MSE optimality of the mean. Therefore, adding $0=\operatorname{E}\left[\,\tfrac{\boldsymbol{\mu}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$ to (6), we find
$\operatorname{E}\left[\,\frac{\mathbf{m}_{p}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]=\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$
Pulling constants out of the expectations and dividing by $\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]>0$ (otherwise $\mathbf{X}=\mathbf{m}_{p}$ almost surely, a contradiction to $\sigma_{\min}>0$), we have shown the claim. ∎
See 2.2
Proof.
By assumption, we have $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ for all $i\in\mathbb{N}$. Using the triangle inequality, optimality of the geometric median, Jensen’s inequality and the variance bound yields
$\|\mathbf{m}_{p}-\boldsymbol{\mu}\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\boldsymbol{\mu}\,\right]\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\mathbf{X}\,\right]\|\leq\operatorname{E}\left[\,\|\mathbf{m}_{p}-\mathbf{X}\|\,\right]\leq\\ \operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|\,\right]\leq\sqrt{\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|^{2}\,\right]}\leq\sqrt{Cp}$
as desired. ∎
See 2.3
Proof.
To simplify notation and stress the additive structure, we denote $S(\mathbf{v}_{p}):=\sum_{i=1}^{p}(Y_{i}-v_{p,i})^{2}=\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}$ and $s_{p}:=\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$ for $p\in\mathbb{N}$. The quantity of interest thus is $\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]$.
Let $k>0$ and $p\geq\lceil 4k\rceil(M+1)+1$. We use $n:=\lfloor\tfrac{p-1}{M+1}\rfloor+1$, and note $n\geq 4k$.
We repeatedly make use of the following bounds on $s_{p}$
$\sigma_{\min}^{2}p\leq s_{p}\leq\sum_{i=1}^{p}\operatorname{E}\left[\,Y_{i}^{2}\,\right]+\operatorname{E}\left[\,\sum_{i=1}^{p}2|Y_{i}\cdot v_{p,i}|\,\right]+\|\mathbf{v}_{p}\|^{2}\leq\\ Cp+\operatorname{E}\left[\,\sum_{i=1}^{p}|Y_{i}|^{2}+(v_{p,i})^{2}\,\right]+C_{2}p\leq 2(C+C_{2})p,$ (7)
where the lower bound follows from $\operatorname{E}\left[\,(Y_{i}-v_{p,i})^{2}\,\right]\geq\operatorname{E}\left[\,(Y_{i}-\operatorname{E}\left[\,Y_{i}\,\right])^{2}\,\right]=\sigma_{\min}^{2}$, and we used Cauchy’s inequality $|ab|\leq\tfrac{a^{2}}{2}+\tfrac{b^{2}}{2}$.
For the first step of the proof, recall that the expectation of a nonnegative random variable $Y$ can be written as $\int_{0}^{\infty}\operatorname{P}\left(\,Y\geq t\,\right)\mathrm{d}t$, irrespective of whether it is finite or not. Thus
$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]=\int_{0}^{\infty}\operatorname{P}\left(\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\geq t\,\right)\mathrm{d}t=\int_{0}^{\infty}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t.$ (8)
We split the integral in (8) into three parts, obtained by splitting the positive reals at the positions $r^{k}\leq R^{k}$, where
$r$ $:=48(C+C_{2})/\sigma_{\min}^{2}$ $R$ $:=r\cdot p^{6}.$ (9)
We bound the integrand in three different ways, depending on the segment.
Step 1: upper most segment Let $t\in(R^{k},\infty)$, meaning we need to bound the probability that $S(\mathbf{v}_{p})$ is close to zero. Since zero has a finite distance to $\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$, common moment-based concentration inequalities result in a constant bound as $t\to\infty$. This is prohibits their application here, since we need an integrable bound. Instead, we use the existence of densities and their uniform boundedness to obtain a bound that vanishes as $t\to\infty$.
Note that if $X$ has density $f$, then, for arbitrary $m\in\mathbb{R}$, the density of $(X-m)^{2}$ is $0.5(f(m+\sqrt{x})+f(m-\sqrt{x}))/\sqrt{x}$ with $x\in\mathbb{R}_{+}$. Consequently, for all $i\in\mathbb{N},m\in\mathbb{R}$ and $\tau\geq 0$,
$\operatorname{P}\left(\,(Y_{i}-m)^{2}\leq\tau\,\right)\leq\int_{0}^{\tau}\frac{c}{\sqrt{x}}\mathrm{d}x=2c\sqrt{\tau}.$ (10)
We aggregate this bound on an independent subset of $(Y_{i})_{i=1,\ldots,p}$. Note that by construction of $n$, it holds that $(n-1)(M+1)+1\leq p$. By $M$-dependence, $(Y_{j(M+1)+1})_{j=0,\ldots,n-1}$ are independent random variables. Thus
[…]
Indeed, this bound is integrable in $t$:
$\int_{R^{k}}^{\infty}t^{-\frac{n}{2k}}\mathrm{d}t=\frac{1}{-\frac{n}{2k}+1}\left[t^{-\frac{n}{2k}+1}\right]_{R^{k}}^{\infty}\overset{\frac{-n}{2k}+1<0}{=}\\ \frac{1}{\frac{n}{2k}-1}R^{-\frac{n}{2}+k}=\frac{1}{\frac{n}{2k}-1}p^{-3n+6k}r^{-\frac{n}{2}+k}\overset{n\geq 4k}{\leq}p^{-3n+6k}r^{-\frac{n}{2}+k}.$
Thus, the contribution of the upper most segment to the expectation of interest is
[…]
Step 2: middle segment Let $t\in(r^{k},R^{k})$. Then,
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\leq\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right).$ (13)
The right hand side probability, i.e. that a sum of random variables deviates from its expectation, can efficiently be treated by concentration inequalities from the literature. Here, we apply a Hoeffding-type bound allowing for partial dependence among the summands. Since Hoeffding requires bounded random variables, we begin by approximating $S(\mathbf{v}_{p})$ and $s_{p}$ by quantities involving bounded random variables.
According to Lemma A.3, for $B:=\max(\sqrt{2C_{2}},\bar{B})$ with some $\bar{B}$ only depending on global constants, it holds that
$\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\frac{\sigma_{\min}^{2}}{2}$
for all $1\leq i\leq p$. We set $\mathcal{I}:=\{1\leq i\leq p\mid|v_{p,i}|\leq B\}$. Note that $\mathcal{I}$ contains at least $\lfloor\tfrac{p}{2}\rfloor$ elements. Otherwise,
$\|\mathbf{v}_{p}\|^{2}\geq\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}v_{p,i}^{2}>\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}(\sqrt{2C_{2}})^{2}\geq\lceil\frac{p}{2}\rceil 2C_{2}\geq C_{2}p,$
a contradiction to the theorem assumptions. Defining
$A_{p,i}$ $:=(\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)-v_{p,i})^{2}$ $a_{p}$ $:=\operatorname{E}\left[\,\sum_{i\in\mathcal{I}}A_{p,i}\,\right],$
we note that by the previous discussion
$a_{p}\geq\sum_{i\in\mathcal{I}}\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\lfloor\tfrac{p}{2}\rfloor\cdot\frac{\sigma_{\min}^{2}}{2}.$ (14)
Note that for $i\in\mathcal{I}$, we have $A_{p,i}\leq(Y_{i}-v_{p,i})^{2}$. Consequently, $S(\mathbf{v}_{p})\geq\sum_{i\in\mathcal{I}}(Y_{i}-v_{p,i})^{2}\geq\sum_{i\in\mathcal{I}}A_{p,i}$ and we have
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq\frac{s_{p}}{r}\,\right)=\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-(a_{p}-\frac{s_{p}}{r})\,\right).$ (15)
We have
[…]
Recalling $r=48(C+C_{2})/\sigma_{\min}^{2}$, we have
[…]
If $p=2$, then also $\tfrac{2(C+C_{2})p}{r}\leq\lfloor\tfrac{p}{2}\rfloor\tfrac{1}{2}\tfrac{\sigma_{\min}^{2}}{2}\leq\tfrac{a_{p}}{2}$. Thus, in particular for the $p$ fixed at the beginning of the proof, by (16) it holds that
$a_{p}-\frac{s_{p}}{r}\geq\frac{a_{p}}{2}\geq\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}.$
Inserting this in (15), we obtain
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right).$ (17)
We now apply the Hoeffding-type bound. First, note that for $i\in\mathcal{I}$, we have $0\leq A_{p,i}\leq(2B)^{2}$ because both $|\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)|\leq B$ and $|v_{p,i}|\leq B$.
The Hoeffding-type bound from [7, Thm. 2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$. By $M$-dependence, we may cover $\{1,\ldots,p\}$ by $M+1$ sets of independent variables: consider the index sets $\mathcal{I}_{j}:=\{j+l(M+1):l=0,\ldots,\big\lfloor\frac{p-j}{M+1}\big\rfloor\}.$ for $j=1,\ldots,M+1$. Then $(A_{p,i})_{i\in\mathcal{I}_{j}}$ are independent for all fixed $j=1,\ldots,M+1$. Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [7]. Thus
$\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right)\leq\exp\left(-2\frac{(\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4})^{2}}{(M+1)\sum_{i\in\mathcal{I}}(2B)^{4}}\right)\leq\\ \exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right).$
Inserting this into equation (17), we can bound the contribution of the middle segment to the expectation of interest as
[…]
Step 3: lower segment Let $t\in(0,r^{k})$. Since $r^{k}$ is finite and does not depend on $p$, it suffices to bound the integrated probability by $1$. Thus, the contribution of the lower segment to the expectation of interest is
$\int_{0}^{r^{k}}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t\leq r^{k}.$ (19)
Final step It remains to aggregate the three segments. Inserting (12), (18), (19) into (8), we have shown
$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]\leq r^{k}\left(\frac{2c\sqrt{2(C+C_{2})}}{p\sqrt{r}}\right)^{n}p^{-3n/2+6k}+\\ r^{k}\cdot\exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right)\cdot p^{6k}+r^{k}.$
We have shown $\sup_{p}\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]<\infty$ if the right hand side remains bounded as $p\to\infty$. Indeed, the basis of the first term is smaller than $1$ if $p$ is large enough and the exponent is $n\geq 4k>0$; further by $n\geq 4k$ we have $3n/2+6k\leq 0$, and finally the exponential decay by $\lfloor\tfrac{p}{2}\rfloor^{2}/p$ is faster than the growth of $p^{6k}$.
Note that the right hand side of the bound directly only depends on $c,C$ and $C_{2}$, $k$, $\sigma_{\min}^{2}$ and $M$ and that indirectly $r,n,B$ are functions of $M,q,p,C,C_{2},\sigma_{\min}^{2}$ only. Thus, the finite supremum of the right hand side also only depends on those quantities. ∎
Lemma A.3.
Let $(Y_{i})_{i\in\mathbb{N}}$ be a sequence of real random variables such that $\sup_{i\in\mathbb{N}}\operatorname{E}\left[\,|Y_{i}|^{q}\,\right]\leq C_{1}<\infty$ for some $q>2$ and $\inf_{i\in\mathbb{N}}\operatorname{Var}\left[\,Y_{i}\,\right]\geq\sigma_{\min}^{2}$. Then there exists $\bar{B}>0$ depending only $q,C_{1}$ and $\sigma_{\min}^{2}$ such for all $B\geq\bar{B}$
$\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]\geq\frac{\sigma_{\min}^{2}}{2}$
for all $p\in\mathbb{N}$ and $1\leq i\leq p$.
Proof.
First, recall that for any nonnegative random variable $Y$ and $j>0$, we have $\operatorname{E}\left[\,Y^{j}\,\right]=\int_{0}^{\infty}jy^{j-1}\operatorname{P}\left(\,Y>y\,\right)\mathrm{d}y$ [4, Lemma 2.2.13].
Let $B>0$ be arbitrary. Then, for $j\in\{1,2\}$, $i,p\in\mathbb{N}$, using Markov’s inequality, we have
$\operatorname{E}\left[\,|\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|^{j}\cdot 1_{\{|Y_{i}|>B\}}\,\right]=B^{j}\operatorname{P}\left(\,|Y_{i}|>B\,\right)=\\ B^{j}\operatorname{P}\left(\,|Y_{i}|^{q}>B^{q}\,\right)\leq B^{j}\frac{C_{1}}{B^{q}}=C_{1}B^{j-q}.$
Further, note $\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\cdot 1_{\{|Y_{i}|\leq B\}}=Y_{i}\cdot 1_{\{|Y_{i}|\leq B\}}$. Thus, we have
$\left|\operatorname{Var}\left[\,Y_{i}\,\right]-\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]\right|=|\operatorname{E}\left[\,Y_{i}^{2}\,\right]-\operatorname{E}\left[\,Y_{i}\,\right]^{2}-\\ \operatorname{E}\left[\,(\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B))^{2}\,\right]+\operatorname{E}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]^{2}|\leq\\ \operatorname{E}\left[\,(\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B))^{2}\cdot 1_{\{|Y_{i}|>B\}}\,\right]+\\ \operatorname{E}\left[\,|\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|\cdot 1_{\{|Y_{i}|>B\}}\,\right]\cdot\operatorname{E}\left[\,|Y_{i}+\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|\,\right]\leq\\ C_{1}B^{2-q}+C_{1}B^{1-q}\cdot(C_{1}^{1/q}+B)=2C_{1}B^{2-q}+C_{1}^{1+1/q}B^{1-q}.$
Letting $B\to\infty$, we obtain that the difference of the variances is less than $\tfrac{\sigma_{\min}^{2}}{2}$ for $B$ large enough, which, together with the fact that the bound in the preceeding display is monotonically decreasing in $B$, yields the claim. ∎
See 2.4
Proof.
First, consider $p\geq 4(M+1)+1$ and $1\leq i\leq p$. Note that by Theorem 2.3, $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{-2}\,\right]$ exists, which makes the expectations in this proof well-defined.
Taking (1) componentwise, we find
$|\mu_{i}-m_{p,i}|\leq\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left|\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]\right|.$ (20)
Applying Jensen’s inequality with the convex function $1/\sqrt{x}$, we have
$\frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\leq\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}=1.$
Further, by the Cauchy-Schwarz inequality
[…]
where we applied Theorem 2.3 with $k=1$, using that $\|\mathbf{m}_{p}\|^{2}\leq Cp$ by Lemma 2.2. Inserting both the previous bounds into equation (20) yields the claim for $p\geq 4(M+1)+1$.
When $p<4(M+1)+1$, use Lemma 2.2 and properties of $\mathcal{D}$ to bound
$|\mu_{i}-m_{p,i}|\leq\|\boldsymbol{\mu}-\mathbf{m}_{p}\|\leq\sqrt{Cp}\leq\sqrt{C\cdot(4(M+1)+1)},$
which finalizes the proof. ∎
The following Lemma generalizes Lemma 2.3 to $\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}$ and generalizations thereof. A challenge for this task is that $(\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2})_{p\in\mathbb{N}}$ is not of the form that Lemma 2.3 gives guarantees for, simply because its number of summands does not grow while $|p-i|\leq M$. One can cut out the stagnating region from the sequence (which we we could have done for the asymptotic case $p\gg i$), but then the sequence never contains $\mathbf{m}_{p}$ for $|p-i|\leq M$, which is a problem if one is interested in all $p$. The following Lemma, for some fixed $p$, constructs a sequence admissible for Theorem 2.3 which, for example, agrees with $\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}$ in one index, but may not equal $(\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2})_{p\in\mathbb{N}}$ in other indices.
Definition A.4.
For $i,n\in\mathbb{N}$, define the exclusion sets $\mathcal{E}^{n}(i):=\{i-nM,\ldots,i+nM\}$. We write $\mathcal{E}(i):=\mathcal{E}^{1}(i)$. For $\mathbf{v}\in\mathbb{R}^{p}$, define $\mathbf{v}_{-\mathcal{E}^{n}(i)}:=(v_{j})_{j\in\{1,\ldots,p\}\setminus\mathcal{E}^{n}(i)}$.
Lemma A.5.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, and $k>0,n\in\mathbb{N}$. When $p\geq\lceil 2k\rceil(M+1)+2(nM+1)$, then for any $1\leq i\leq p$
$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{k}}\,\right]\leq\frac{B_{2}(k,n,\mathcal{D})}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}}$ (21)
for a constant $B_{2}<\infty$ only depending on $k,n$ and the $\mathcal{D}$-parameters.
Proof.
Let $p\geq\lceil 2k\rceil(M+1)+2(nM+1)$ and $1\leq i\leq p$ be fixed. Define the exclusion set $\mathcal{E}:=\mathcal{E}^{n}(i)\cap\{1,\ldots,p\}$ and the index set $\check{\mathcal{J}}:=\{1,\ldots,p\}\setminus\mathcal{E}$. By definition, $\mathbf{X}_{-\mathcal{E}^{n}(i)}=(X_{j})_{j\in\check{\mathcal{J}}}$ and $\mathbf{m}_{p,-\mathcal{E}^{n}(i)}=(m_{p,j})_{j\in\check{\mathcal{J}}}$.
Let $\mathcal{J}:=\mathbb{N}\setminus\mathcal{E}$. With $l\in\mathbb{N}$ denoting the sequence index, we plan to apply Theorem 2.3 to $Y_{l}:=X_{\mathcal{J}_{l}}$, where $\mathcal{J}_{l}$ denotes the $l$-th entry of $\mathcal{J}$, and to $\mathbf{v}_{l}:=(m_{l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}\,,\,\mathcal{J}_{j}})_{j=1,\ldots,l}\in\mathbb{R}^{l}$, where $|\cdot|$ denotes set cardinalities and $x_{+}=\max(x,0)$ for $x\in\mathbb{R}$. To see that $\mathbf{v}_{l}$ is well-defined (recall that $m_{a,b}$ requires $b\leq a$), we check $\mathcal{J}_{j}\leq\mathcal{J}_{l}\leq l+|\mathcal{E}|\leq l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}$ for all $1\leq j\leq l$. Regarding the remaining theorem assumptions, note that, just like $(X_{l})_{l}$, we have $(Y_{l})_{l}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$. Further,
[…]
where in the final step we used $1\leq l$.
Thus, applying Theorem 2.3 for $k/2$, with $s_{l}:=\operatorname{E}\left[\,\sum_{j=1}^{l}(Y_{j}-v_{l,j})^{2}\,\right]$, we have
$\sup_{l\geq\lceil 2k\rceil(M+1)+1}\operatorname{E}\left[\,\left(\frac{1}{s_{l}}\sum_{j=1}^{l}(Y_{j}-v_{l,j})^{2}\right)^{-k/2}\,\right]\leq\beta(k/2,2C+C(2nM+1),\mathcal{D})<\infty.$
For translating this result to $\mathbf{X}_{-\mathcal{E}^{n}(i)}$, note that when setting $l:=|\check{\mathcal{J}}|$, we have $(Y_{1},\ldots,Y_{|\check{\mathcal{J}}|})=\mathbf{X}_{-\mathcal{E}^{n}(i)}$ and $\mathbf{v}_{l}=\mathbf{m}_{p,-\mathcal{E}^{n}(i)}$ by $l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}=|\check{\mathcal{J}}|+|\mathcal{E}|+0=p$. Further, using the assumption $p\geq(\lceil 2k\rceil+2n))(M+1)$, we have
$|\check{\mathcal{J}}|=p-|\mathcal{E}|\geq p-(2nM+1)\geq\\ \lceil 2k\rceil(M+1)+2(nM+1)-(2nM+1)=\lceil 2k\rceil(M+1)+1,$
which ensures the sup-constraint is satisfied for $l:=|\check{\mathcal{J}}|$. Thus, in particular we have
$\operatorname{E}\left[\,\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}}{\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{k}}\,\right]\leq\beta(k/2,2C+C(2nM+1),\mathcal{D})<\infty,$
which, after dividing by $\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}\geq(|\check{\mathcal{J}}|\sigma_{\min}^{2})^{k/2}>0$, yields the claim. ∎
Lemma A.6.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$ and $0\leq A,B<\infty$ be constants. Then, for any $i,j\in\mathbb{N}$ we have
$\operatorname{E}\left[\,((|X_{i}|+A)(|X_{j}|+B)^{2})^{q/3}\,\right]\leq B_{4}(A,B,\mathcal{D})<\infty$ (22)
for a constant $B_{4}$ only depending on $A,B$ and the $\mathcal{D}$-parameters
Proof.
First, note that
$(|X_{i}|+A)(|X_{j}|+B)^{2}=|X_{i}|X_{j}^{2}+2|X_{i}||X_{j}|B+|X_{i}|B^{2}+AX_{j}^{2}+2A|X_{j}|B+AB^{2}.$
Noting that by $q\geq 3$ the function $x\mapsto x^{q/3}$ is convex, we have by the convexity definition that
$((|X_{i}|+A)(|X_{j}|+B)^{2})^{q/3}\leq\frac{6^{q/3}}{6}\bigg((|X_{i}|X_{j}^{2})^{q/3}+(2|X_{i}||X_{j}|B)^{q/3}+\\ (|X_{i}|B^{2})^{q/3}+(AX_{j}^{2})^{q/3}+(2A|X_{j}|B)^{q/3}+(AB^{2})^{q/3}\bigg).$ (23)
By Jensen’s inequality, the definition of the $\mathcal{D}$-class (in particular $C\geq 1$), it holds that $\operatorname{E}\left[\,(|X_{i}|^{k})^{q/3}\,\right]\leq\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{k/3}\leq C^{k/3}\leq C$ for all $0<k\leq 2$. The same holds for $|X_{j}|$.
By the Cauchy-Schwarz inequality and Jensen’s inequality, it holds that
$\operatorname{E}\left[\,(|X_{i}||X_{j}|)^{q/3}\,\right]\leq\sqrt{\operatorname{E}\left[\,|X_{i}|^{2q/3}\,\right]}\sqrt{\operatorname{E}\left[\,|X_{j}|^{2q/3}\,\right]}\leq\\ \sqrt{\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{2/3}}\sqrt{\operatorname{E}\left[\,|X_{j}|^{q}\,\right]^{2/3}}\leq C^{1/3}C^{1/3}\leq C.$
Finally, applying Hoelder’s inequality, we find
$\operatorname{E}\left[\,(|X_{i}|X_{j}^{2})^{q/3}\,\right]\leq\||X_{i}|^{q/3}\|_{L^{3}}\||X_{j}|^{2q/3}\|_{L^{3/2}}=\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{1/3}\operatorname{E}\left[\,|X_{j}|^{q}\,\right]^{2/3}\leq C.$
Taking the expectation of (23) and inserting the previous three inequalities yields the claim. ∎
See 2.5
Proof.
First consider $p\geq 14M+8$. Let $1\leq i\leq p$ be fixed and set $\mathcal{E}:=\mathcal{E}^{1}(i)\cap\{1,\ldots,p\}$. Recall $\check{\mathbf{X}}=(X_{i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}}$ and $\check{\mathbf{m}}_{p}=(m_{p,i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}}$. By $M$-dependence of $(X_{i})_{i}$, we have that
$\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]=\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\operatorname{E}\left[\,\frac{1}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]=0,$
since $\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]=0$ and the second factor is finite by Lemma A.5. Thus,
$|m_{p,i}-\mu_{i}|=\left|\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left(\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]-\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]\right)\right|=\\ \left|\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]\right|\leq\\ \operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{|(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}|}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\,\right]=\\ \operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\,\right]\operatorname{E}\left[\,|(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}|\,\right]\leq\\ \frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\cdot\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]^{3/2}}\cdot\\ \operatorname{E}\left[\,\left(\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\right)^{3/2}\,\right]\cdot\sum_{j\in\mathcal{E}}\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|\,\right]\\ =:I\cdot II\cdot III\cdot IV,$
where we used $\sqrt{x}-\sqrt{y}=\tfrac{x-y}{\sqrt{x}+\sqrt{y}}$, followed by the triangle inequality and $\|\mathbf{X}-\mathbf{m}_{p}\|\geq\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\geq\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|\geq 0$, and finally independence of $(X_{j})_{j\in\mathcal{E}}$ and $(X_{j})_{j\in\{1,\ldots,p\}\setminus\mathcal{E}^{2}(i)}$ because of $M$-independence.
For the first Term $I$, it holds that $I\leq 1$ using Jensen’s inequality as shown in the proof of Lemma 2.4.
Regarding Term $II$, note that
$\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}\,\right])+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2},$
since $\operatorname{E}\left[\,(\mathbf{X}-\boldsymbol{\mu})^{T}(\boldsymbol{\mu}-\mathbf{m}_{p})\,\right]=0$, and similarly $\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}_{-\mathcal{E}^{2}(i)}\,\right])+\|\boldsymbol{\mu}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|$. Using Lemma 2.2, properties of $\mathcal{D}$, and $4M+1\leq\tfrac{1}{3}p$ by $p\geq 14M+8$, we have
$II\leq\frac{\sqrt{Cp+Cp}}{((p-(4M+1))\sigma_{\min}^{2}+0)^{3/2}}\leq\frac{\sqrt{p}\sqrt{2C}}{(\tfrac{2}{3}p)^{3/2}\,\sigma_{\min}^{3}}=\frac{1}{p}\cdot\frac{\sqrt{2C}}{\sigma_{\min}^{3}\cdot(\tfrac{2}{3})^{3/2}}.$
Applying Lemma A.5 for $k:=3$ and $n:=2$ to Term $III$, we have $III\leq B_{2}(3,2,\mathcal{D})$.
For Term $IV$, first note $|\mu_{i}|\leq C$ and $|m_{p,j}|\leq B_{1}+C$ according to Lemma 2.4. Thus, applying Jensen’s inequality (note $x\mapsto x^{3/q}$ is concave by $q\geq 3$) and Lemma A.6, we find that for any $j\in\mathcal{E}$ we have
$\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|\,\right]\leq\operatorname{E}\left[\,(|X_{i}|+C)(|X_{j}|+B_{1}+C)^{2}\,\right]\leq\\ \operatorname{E}\left[\,((|X_{i}|+C)(|X_{j}|+B_{1}+C)^{2})^{q/3}\,\right]^{3/q}\leq B_{4}(C,B_{1}+C,\mathcal{D})^{3/q}.$
Since $\mathcal{E}$ contains at most $2M+1$ indices, we have shown that term $IV$ is bounded with constant only depending on the $\mathcal{D}$-parameters.
In summary, Terms $I,III$ and $IV$ are bounded and Term $II$ is $\mathcal{O}(\tfrac{1}{p})$, which yields the Theorem claim for $p\geq 14M+8$.
Let $p\leq 14M+8$. Then, for all $1\leq i\leq p$, by Lemma 2.2,
$|m_{p,i}-\mu_{i}|\leq\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\sqrt{Cp}=\frac{1}{p}\sqrt{C}p^{3/2}\leq\frac{1}{p}\sqrt{C}(14M+8)^{3/2}.$
Taking the maximum of the constants derived in the two cases for $p$ yields the theorem claim. ∎
See 2.6
Proof.
We show
$p\cdot(m_{p,i}-\mu_{i})\xrightarrow{p\to\infty}\frac{1}{2\bar{\sigma}}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right],$
which is equivalent to the claim. Similar to the proof of Theorem 2.5, one can see that for $p\geq i+M$
$p\cdot(m_{p,i}-\mu_{i})=p\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left(\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]-\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]\right)=\\ p\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]=\\ \frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\quad\cdot\quad\frac{p\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{3/2}}\quad\cdot\\ \operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{3/2}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]\\ =:I\cdot II\cdot III.$ (24)
Starting with the first Term, $I$, we note that it can be rewritten as
$I=\operatorname{E}\left[\,\left(\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\right)^{-1/2}\,\right]^{-1}.$
Since by Lemma A.7 we have $\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1$ in probability as $p\to\infty$, also $(\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]})^{-1/2}\to 1$ by the continuos mapping theorem. To show that the random sequence’s expectations also converge, we appeal to a uniform integrability argument [3, Thm 6.2]. Indeed, by Theorem 2.3 it holds for example that
$\sup_{p\geq 4(M+1)+1}\operatorname{E}\left[\,\left(\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\right)^{-1}\,\right]\leq\beta(1,C,\mathcal{D})<\infty.$
[…]
Regarding Term $II$, recall from the proof of Theorem 2.5 that $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}\,\right])+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}$. Further, by Theorem 2.5 it holds that $\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}=\mathcal{O}(\tfrac{1}{p})$ for $p$ large enough. Consequently, $\tfrac{1}{p}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\to\bar{\sigma}^{2}$. Since $\bar{\sigma}\geq\sigma_{\min}>0$, the second Term converges to $\tfrac{1}{\bar{\sigma}^{2}}$.
We treat Term $III$ similar to Term $I$, i.e. first show convergence in probability of the expectation argument and then appeal to a uniform integrability argument to deduce convergence of the expectations. To simplify notation, rewrite $III=\operatorname{E}\left[\,N_{p}/D_{p}\,\right]$, where the numerator $N_{p}$ and denominator $D_{p}$ are given by
$N_{p}$ $:=-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2},$
$D_{p}$ $:=\sqrt{\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\cdot\sqrt{\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\left(\sqrt{\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}+\sqrt{\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\right).$
Starting with the numerator $N_{p}$, note that $m_{p,j}\to\mu_{j}$ by Theorem 2.5. Since $\mathcal{E}(i)$ contains finitely many terms, we conclude by the continuous mapping theorem that $N_{p}\to N:=-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-\mu_{j})^{2}$ in probability (note that for convergence of a random vector in probability it suffices that the components converge in probability).
For studying the denominator $D_{p}$, note that for $p\geq i+M$,
$\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}=\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}-\frac{\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\xrightarrow{p\to\infty}1-0=1$
in probability, since the first term converges to $1$ by Lemma A.7, and the numerator of the second term converges to $\sum_{j\in\mathcal{E}(i)}(X_{j}-\mu_{j})^{2}$ as discussed while $\tfrac{1}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\leq\tfrac{1}{p\sigma_{\min}^{2}}\to 0$, implying the second fraction converges to 0 by Slutsky’s theorem.
Thus, another application of the continuous mapping theorem implies $D_{p}\to 2$ in probability. All in all, $N_{p}/D_{p}\to N/2$ in probability by Slutsky’s theorem, where $N$ is as defined previously.
Next, we show $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$. Crucially $q/3>1$, which is required for the uniform integrability argument. First, note that similar to the proof of Theorem 2.5,
$\operatorname{E}\left[\,\left(\frac{|N_{p}|}{D_{p}}\right)^{q/3}\,\right]\leq\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}\operatorname{E}\left[\,\left(\frac{|N_{p}|}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\right)^{q/3}\,\right]=\\ \frac{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]^{q/2}}\cdot\operatorname{E}\left[\,|N_{p}|^{q/3}\,\right]\cdot\\ \operatorname{E}\left[\,\left(\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\right)^{q/2}\,\right]=:III^{\prime}\cdot III^{\prime\prime}\cdot III^{\prime\prime\prime}.$
We have, using Lemma 2.2, that for $p\geq 4M+2$
$(III^{\prime})^{2/q}=\frac{\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}\,\right]}+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}}{\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}_{-\mathcal{E}^{2}(i)}\,\right]}+\|\boldsymbol{\mu}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\leq\\ \frac{Cp+Cp}{(p-|\mathcal{E}^{2}(i)|)\sigma_{\min}^{2}+0}\leq\frac{2Cp}{(p-(4M+1))\sigma_{\min}^{2}},$
which remains bounded as $p\to\infty$ with bound only depending on the $\mathcal{D}$-parameters.
For the second term, use convexity of $x\mapsto x^{q/3}$, the fact that $|\mu_{i}|\leq C$ and $|m_{p,j}|\leq B_{1}+C$, and Lemma A.6 to find
$III^{\prime\prime}\leq|\mathcal{E}(i)|^{q/3}\sum_{j\in\mathcal{E}(i)}\frac{1}{|\mathcal{E}(i)|}\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|^{q/3}\,\right]\leq\\ |\mathcal{E}(i)|^{(q-3)/3}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,((|X_{1}|+C)(|X_{j}|+B_{1}+C)^{2})^{q/3}\,\right]\leq\\ (2M+1)^{(q-3)/3}(2M+1)B_{4}(C,B_{1}+C,\mathcal{D}).$
Finally, when $p$ large enough, then by Lemma A.5, we have $III^{\prime\prime\prime}\leq B_{2}(q,2,\mathcal{D})$.
In summary, $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$ when $p$ large enough. By [3, Thm 6.2], we conclude $III\to\operatorname{E}\left[\,N/2\,\right]$ as $p\to\infty$.
Inserting the obtained limits for $I$, IIII, and $III$ into (24) yields the claim. ∎
Lemma A.7.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Then, $\|\mathbf{X}-\mathbf{m}_{p}\|^{2}/\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\to 1$ in probability.
Proof.
If $q\geq 4$, the claim can be shown using Chebychev’s inequality. To allow $q>3$ in Theorem 2.6, we instead employ a truncation approach similar to [4, Sect. 2.2.3]. For $1\leq i\leq p$, define $A_{p,i}:=(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}$ and set $a_{p}:=\operatorname{E}\left[\,\sum_{i=1}^{p}A_{p,i}\,\right]$.
First, note by the triangle inequality, convexity of $x\mapsto x^{q}$, and Lemma 2.4 that
$\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq 2^{q}(\tfrac{1}{2}\operatorname{E}\left[\,|X_{i}|^{q}\,\right]+\tfrac{1}{2}|m_{p,i}|^{q})\leq 2^{q-1}(C+B_{1}^{q}):=\tilde{C}$ (25)
for $1\leq i\leq p$. Further, for $t>0$ by Markov’s inequality,
$\operatorname{P}\left(\,(X_{i}-m_{p,i})^{2}>t\,\right)=\operatorname{P}\left(\,|X_{i}-m_{p,i}|^{q}>t^{q/2}\,\right)\leq\frac{\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]}{t^{q/2}}\leq\frac{\tilde{C}}{t^{q/2}}.$ (26)
Third, using $q>2$, we make use of the bound
[…]
To prove convergence in probability, let $t>0$. Then,
$\operatorname{P}\left(\,\left|\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{a_{p}}-1\right|>t\,\right)=\operatorname{P}\left(\,\left|\|\mathbf{X}-\mathbf{m}_{p}\|^{2}-a_{p}\right|>a_{p}t\,\right)\leq\\ \operatorname{P}\left(\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\neq\sum_{i=1}^{p}A_{p,i}\,\right)+\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right).$ (28)
We have by inclusion of events that by $q>2$,
[…]
We treat the second term with Chebycheff’s inequality. First, when $q\geq 4$, $\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\operatorname{E}\left[\,(X_{i}-m_{p,i})^{4}\,\right]\leq\tilde{C}^{4/q}\leq\tilde{C}$ since $\tilde{C}\geq C\geq 1$ and using Jensen’s inequality. If $q<4$,
$\operatorname{E}\left[\,A_{p,i}^{2}\,\right]=\operatorname{E}\left[\,((X_{i}-m_{p,i})^{2})^{2-q/2}|X_{i}-m_{p,i}|^{q}1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}\,\right]\leq\\ p^{2-q/2}\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq p^{2-q/2}\tilde{C}$.
Thus, in any case, $\operatorname{Var}\left[\,A_{p,i}\,\right]\leq\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\tilde{C}\cdot\max(1,p^{2-q/2})$.
Let $i\in\mathbb{N}$. By $M$-dependence, $\operatorname{Cov}(A_{p,i},A_{p,j})=0$ for all but $2M+1$ indices $j$. This fact together with $|\operatorname{Cov}(X,Y)|\leq\sqrt{\operatorname{Var}\left[\,X\,\right]\operatorname{Var}\left[\,Y\,\right]}$ gives
$\left|\sum_{j=1}^{p}\operatorname{Cov}(A_{p,i},A_{p,j})\right|\leq(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2}).$
Thus,
$\operatorname{E}\left[\,\left(\sum_{i=1}^{p}A_{p,i}-a_{p}\right)^{2}\,\right]=\operatorname{Var}\left[\,\sum_{i=1}^{p}A_{p,i}\,\right]\leq p\cdot(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2})$
Hence, by Chebycheff’s inequality,
$\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right)\leq\frac{p\cdot(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2})}{t^{2}a_{p}^{2}}.$
Note
[…]
Thus, for $p$ large enough,
$\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right)\leq\frac{1}{t^{2}}\cdot\frac{4(2M+1)\tilde{C}}{\sigma_{\min}^{4}}\cdot\max\left(\frac{1}{p},p^{1-q/2}\right)\xrightarrow{p\to\infty}0.$ (30)
Combining (30) with (29) in (28), we have shown
$\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{a_{p}}\xrightarrow{P}1$ (31)
in probability as $p\to\infty$. Further, note
$\frac{\sum_{i=1}^{p}\operatorname{E}\left[\,(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}>p\}}\,\right]}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\leq\frac{p^{2-q/2}}{p\,\sigma_{\min}^{2}}\left(\tilde{C}+\frac{\tilde{C}}{q/2-1}\right)\xrightarrow{p\to\infty}0,$
which implies
$\frac{a_{p}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}=1-\frac{\sum_{i=1}^{p}\operatorname{E}\left[\,(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}>p\}}\,\right]}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\xrightarrow{p\to\infty}1.$
Combining this with (31) yields the claim by Slutsky’s theorem. ∎
A.2 Further analysis of the simulation results
Linear Regression slopes for the simulation curves can be found in Table 1. The Pareto(3.1) results are analyzed in more detail in Figure 2.
Exp[0.75, 1.25] MA(2) Pareto(3.1) Pareto[7,8]
Slope of $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ 0.49-0.49 0.51-0.51 0.28-0.28 0.45-0.45
Slope of Residual 1.45-1.45 1.64-1.64 1.03-1.03 1.67-1.67
Table 1: Linear regression slopes to the simulation curves after $p\geq 100$
Figure 2: A closer look at the Pareto(3.1) simulations. Left: $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ (same as left side of Figure 1) in pink versus $f(p):=\sqrt{p}\cdot|\tfrac{-19.52}{p}+\tfrac{20.16}{p^{1.03}}|$ in black. Indeed, $f$ explains the decay of $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ well. Right: Plot $f(p)$ from 100100 to $10^{5}$ in black and compare with $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ in blue. Even for $p$ so large, $f$ is still not fully parallel to $\mathcal{O}(\tfrac{1}{\sqrt{p}})$.